An office has some 20¢ stamps and some 28¢ stamps. Altogether the office has 210 stamps for a total value of $47.60. How many of each type of stamp does the office have?

Answer by ikleyn(52782)

.
An office has some 20¢ stamps and some 28¢ stamps. Altogether the office has 210 stamps for a total value of $47.60.
How many of each type of stamp does the office have?
~~~~~~~~~~~~~~~~~~~~~~

Solution 1. One equation in one unknown

Let x be the number of the 20c stamps.
Then the number of the 28c stamps is (210-x).

A "value" equation is

20x + 28*(210-x) = 4760.       (1)   (4760 = $47.60 )

Simplify and solve for x.

Solution 2. System of two equations in two unknowns

  x +   y =  210,          (1)
20x + 28y = 4760.          (2)

From (1) express y = 210-x and substitute it into (2). You will get and equation

20x + 28*(210-x) = 4760.   (3)

It is identical to the equation (1) in the Solution 1 above.
Simplify it and solve for x.

After finding x calculate y.